# Composite Flow Matching for Reinforcement Learning with Shifted-Dynamics Data

### v2

@@ Appendix D Algorithm of CompFlow built on Soft-Actor-Critic
We have access to the offline source data $\mathcal{M}_{\text{src}}$ and the empirical behavioral policy $\pi_{\mathcal{D}_{\text{src}}}$, so we bound the performance between $J_{\mathcal{M}_{\text{tar}}}(\pi)$ and $J_{\mathcal{M}_{\text{src}}}(\pi_{\mathcal{D}_{\text{src}}})$. We have
$\eta_{\mathcal{M}_{\text{tar}}}(\pi)-\eta_{\pi_{\mathcal{D}_{\text{src}}}}(\pi)=\underbrace{\eta_{\mathcal{M}_{\text{tar}}}(\pi)-\eta_{\mathcal{M}_{\text{src}}}(\pi_{\mathcal{D}_{\text{src}}})}_{(a)}+\underbrace{\eta_{\pi_{\mathcal{D}_{\text{src}}}}(\pi_{\mathcal{D}_{\text{src}}})-\eta_{\pi_{\mathcal{D}_{\text{src}}}}(\pi)}_{(b)}.$ (9)
To bound (a) term, we use Lemma  E.1 in the second equality
$\eta_{\mathcal{M}_{\text{tar}}}(\pi)-\eta_{\mathcal{M}_{\text{src}}}(\mathcal{D}_{\text{src}})$
$=-\left(\eta_{\mathcal{M}_{\text{src}}}(\mathcal{D}_{\text{src}})-\eta_{\mathcal{M}_{\text{tar}}}(\pi)\right)$
$=-\frac{1}{1-\gamma}\,\mathbb{E}_{(s,a)\sim\rho_{\mathcal{M}_{\text{src}}}^{\mathcal{D}_{\text{src}}}}\left[\mathbb{E}_{s^{\prime}_{\text{src}}\sim p_{\mathcal{M}_{\text{src}}},\,a^{\prime}\sim\pi_{\mathcal{D}_{\text{src}}}}\left[Q^{\pi}_{\mathcal{M}_{\text{tar}}}(s^{\prime}_{\text{src}},a^{\prime})\right]-\mathbb{E}_{s^{\prime}_{\text{tar}}\sim p_{\mathcal{M}_{\text{tar}}},\,a^{\prime}\sim\pi}\left[Q^{\pi}_{\mathcal{M}_{\text{tar}}}(s^{\prime}_{\text{tar}},a^{\prime})\right]\right]$
$=-\frac{1}{1-\gamma}\,\mathbb{E}_{(s,a)\sim\rho_{\mathcal{M}_{\text{src}}}^{\mathcal{D}_{\text{src}}}}\left[\right.$
$\quad\underbrace{\left(\mathbb{E}_{s^{\prime}_{\text{src}}\sim p_{\mathcal{M}_{\text{src}}},\,a^{\prime}\sim\pi_{\mathcal{D}_{\text{src}}}}\left[Q^{\pi}_{\mathcal{M}_{\text{tar}}}(s^{\prime}_{\text{src}},a^{\prime})\right]-\mathbb{E}_{s^{\prime}_{\text{src}}\sim p_{\mathcal{M}_{\text{src}}},\,a^{\prime}\sim\pi}\left[Q^{\pi}_{\mathcal{M}_{\text{tar}}}(s^{\prime}_{\text{src}},a^{\prime})\right]\right)}_{\text{(c)}}$
@@
$\mathbb{E}_{s^{\prime}_{\text{src}}\sim p_{\mathcal{M}_{\text{src}}},\,a^{\prime}\sim\pi_{\mathcal{D}_{\text{src}}}}\left[Q^{\pi}_{\mathcal{M}_{\text{tar}}}(s^{\prime}_{\text{src}},a^{\prime})\right]-\mathbb{E}_{s^{\prime}_{\text{src}}\sim p_{\mathcal{M}_{\text{src}}},\,a^{\prime}\sim\pi}\left[Q^{\pi}_{\mathcal{M}_{\text{tar}}}(s^{\prime}_{\text{src}},a^{\prime})\right]$
$\leq\mathbb{E}_{s^{\prime}_{\text{src}}\sim p_{\mathcal{M}_{\text{src}}}}\left[\sum_{a^{\prime}\in\mathcal{A}}\left|\mathcal{D}_{\text{src}}(a^{\prime}\mid s^{\prime}_{\text{src}})-\pi(a^{\prime}\mid s^{\prime}_{\text{src}})\right|\cdot\left|Q^{\pi}_{\mathcal{M}_{\text{tar}}}(s^{\prime}_{\text{src}},a^{\prime})\right|\right]$
@@
$\leq\frac{2r_{\max}}{1-\gamma}\,\mathbb{E}_{s^{\prime}_{\text{src}}\sim p_{\mathcal{M}_{\text{src}}}}\left[D_{\text{TV}}\left(\pi_{\mathcal{D}_{\text{src}}}(\cdot\mid s^{\prime}_{\text{src}})\,\|\,\pi(\cdot\mid s^{\prime}_{\text{src}})\right)\right],$
@@
where the last inequality comes from the fact that $|Q^{\pi}_{\mathcal{M}_{\text{tar}}}(s^{\prime}_{\text{src}},a^{\prime})|\leq\frac{r_{max}}{1-\gamma}$ and the definition of TV distance.
@@
$(d)$ $=\mathbb{E}_{s^{\prime}\sim p_{\mathcal{M}_{\text{src}}},\,a^{\prime}\sim\pi}\left[Q^{\pi}_{\mathcal{M}_{\text{tar}}}(s^{\prime},a^{\prime})\right]-\mathbb{E}_{s^{\prime}\sim p_{\mathcal{M}_{\text{tar}}},\,a^{\prime}\sim\pi}\left[Q^{\pi}_{\mathcal{M}_{\text{tar}}}(s^{\prime},a^{\prime})\right]$
@@
$\leq\frac{r_{\max}}{1-\gamma}\int_{\mathcal{S}}\left|p_{\mathcal{M}_{\text{src}}}(s^{\prime}\mid s,a)-p_{\mathcal{M}_{\text{tar}}}(s^{\prime}\mid s,a)\right|ds^{\prime}$
$=\frac{2r_{\max}}{1-\gamma}\left[D_{\text{TV}}\left(p_{\mathcal{M}_{\text{src}}}(\cdot\mid s,a)\,\|\,p_{\mathcal{M}_{\text{tar}}}(\cdot\mid s,a)\right)\right],$
where the inequality comes from the fact that $Q^{\pi}_{\mathcal{M}_{\text{tar}}}(s^{\prime},a^{\prime})$ weighted by probability is bounded by $\frac{r_{max}}{1-\gamma}$, the upper bound for all Q-values.
Combine the bounds for (c) and (d), we obtain a bound for the (a) term:
$\eta_{\mathcal{M}_{\text{tar}}}(\pi)-\eta_{\mathcal{M}_{\text{src}}}(\pi_{\mathcal{D}_{\text{src}}})$ $\geq-\frac{2r_{\max}}{(1-\gamma)^{2}}\,\mathbb{E}_{(s,a)\sim\rho^{\pi_{\mathcal{D}_{\text{src}}}}_{\mathcal{M}_{\text{src}}},\,s^{\prime}\sim p_{\mathcal{M}_{\text{src}}}}\left[D_{\text{TV}}\left(\pi_{\mathcal{D}_{\text{src}}}(\cdot\mid s^{\prime})\,\|\,\pi(\cdot\mid s^{\prime})\right)\right]$
$\quad-\frac{2r_{\max}}{(1-\gamma)^{2}}\,\mathbb{E}_{(s,a)\sim\rho^{\pi_{\mathcal{D}_{\text{src}}}}_{\mathcal{M}_{\text{src}}}}\left[D_{\text{TV}}\left(p_{\mathcal{M}_{\text{src}}}(\cdot\mid s,a)\,\|\,p_{\mathcal{M}_{\text{tar}}}(\cdot\mid s,a)\right)\right].$
Now we try to bound term (b).
$\eta_{\mathcal{M}_{\text{src}}}(\pi_{\mathcal{D}_{\text{src}}})-\eta_{\mathcal{M}_{\text{src}}}(\pi)$
$=\frac{1}{1-\gamma}\,\mathbb{E}_{(s,a)\sim\rho^{\pi_{\mathcal{D}_{\text{src}}}}_{\mathcal{M}_{\text{src}}},\,s^{\prime}\sim p_{\mathcal{M}_{\text{src}}}(\cdot\mid s,a)}\left[\mathbb{E}_{a^{\prime}\sim\pi_{\mathcal{D}_{\text{src}}}}\left[Q^{\pi}_{\mathcal{M}_{\text{src}}}(s^{\prime},a^{\prime})\right]-\mathbb{E}_{a^{\prime}\sim\pi}\left[Q^{\pi}_{\mathcal{M}_{\text{src}}}(s^{\prime},a^{\prime})\right]\right]$
$\geq-\frac{1}{1-\gamma}\,\mathbb{E}_{(s,a)\sim\rho^{\pi_{\mathcal{D}_{\text{src}}}}_{\mathcal{M}_{\text{src}}},\,s^{\prime}\sim p_{\mathcal{M}_{\text{src}}}(\cdot\mid s,a)}\left|\mathbb{E}_{a^{\prime}\sim\pi_{\mathcal{D}_{\text{src}}}}\left[Q^{\pi}_{\mathcal{M}_{\text{src}}}(s^{\prime},a^{\prime})\right]-\mathbb{E}_{a^{\prime}\sim\pi}\left[Q^{\pi}_{\mathcal{M}_{\text{src}}}(s^{\prime},a^{\prime})\right]\right|$
$\geq-\frac{1}{1-\gamma}\,\mathbb{E}_{(s,a)\sim\rho^{\pi_{\mathcal{D}_{\text{src}}}}_{\mathcal{M}_{\text{src}}},\,s^{\prime}\sim p_{\mathcal{M}_{\text{src}}}(\cdot\mid s,a)}\left|\sum_{a^{\prime}\in\mathcal{A}}\left(\pi_{\mathcal{D}_{\text{src}}}(a^{\prime}\mid s^{\prime})-\pi(a^{\prime}\mid s^{\prime})\right)Q^{\pi}_{\mathcal{M}_{\text{src}}}(s^{\prime},a^{\prime})\right|$
$\geq-\frac{r_{\max}}{(1-\gamma)^{2}}\,\mathbb{E}_{(s,a)\sim\rho^{\pi_{\mathcal{D}_{\text{src}}}}_{\mathcal{M}_{\text{src}}},\,s^{\prime}\sim p_{\mathcal{M}_{\text{src}}}(\cdot\mid s,a)}\left|\sum_{a^{\prime}\in\mathcal{A}}\left(\pi_{\mathcal{D}_{\text{src}}}(a^{\prime}\mid s^{\prime})-\pi(a^{\prime}\mid s^{\prime})\right)\right|$
$=-\frac{2r_{\max}}{(1-\gamma)^{2}}\,\mathbb{E}_{(s,a)\sim\rho^{\pi_{\mathcal{D}_{\text{src}}}}_{\mathcal{M}_{\text{src}}},\,s^{\prime}\sim p_{\mathcal{M}_{\text{src}}}(\cdot\mid s,a)}\left[D_{\text{TV}}\left(\pi_{\mathcal{D}_{\text{src}}}(\cdot\mid s^{\prime})\,\|\,\pi(\cdot\mid s^{\prime})\right)\right],$
where the first equality is a direct application of Lemma  E.2. Combine the bound for term (a) and (b), we conclude that
$\eta_{\mathcal{M}_{\text{tar}}}(\pi)-\eta_{\mathcal{M}_{\text{src}}}(\pi)\geq\;$ $-2C_{1}\,\mathbb{E}_{(s,a)\sim\rho_{\mathcal{M}}^{\pi_{\mathcal{D}_{\text{src}}}},s^{\prime}\sim p_{\text{src}}}\left[D_{\mathrm{TV}}(\pi(\cdot|s^{\prime})\parallel\pi_{\mathcal{D}_{\text{src}}}(\cdot|s^{\prime}))\right]$
$-C_{1}\,\mathbb{E}_{(s,a)\sim\rho_{\mathcal{M}}^{\pi_{\mathcal{D}_{\text{src}}}}}\left[D_{\mathrm{TV}}(p_{\text{tar}}(\cdot|s,a)\parallel p_{\text{src}}(\cdot|s,a))\right],$
where $C_{1}=\frac{2r_{\max}}{(1-\gamma)^{2}}$. ∎
Appendix F Proof of Theorem 3.1

### v3

@@ Appendix E Algorithm of CompFlow built on Soft-Actor-Critic
We have access to the offline data $\mathcal{M}_{\text{off}}$ and the empirical behavioral policy $\pi_{\mathcal{D}_{\text{off}}}$, so we bound the performance between $\eta_{\mathcal{M}_{\text{on}}}(\pi)$ and $\eta_{\mathcal{M}_{\text{off}}}(\pi_{\mathcal{D}_{\text{off}}})$. We have
$\eta_{\mathcal{M}_{\text{on}}}(\pi)-\eta_{\pi_{\mathcal{D}_{\text{off}}}}(\pi)=\underbrace{\eta_{\mathcal{M}_{\text{on}}}(\pi)-\eta_{\mathcal{M}_{\text{off}}}(\pi_{\mathcal{D}_{\text{off}}})}_{(a)}+\underbrace{\eta_{\pi_{\mathcal{D}_{\text{off}}}}(\pi_{\mathcal{D}_{\text{off}}})-\eta_{\pi_{\mathcal{D}_{\text{off}}}}(\pi)}_{(b)}.$ (9)
To bound (a) term, we use Lemma  F.1 in the second equality
$\eta_{\mathcal{M}_{\text{on}}}(\pi)-\eta_{\mathcal{M}_{\text{off}}}(\mathcal{D}_{\text{off}})$
$=-\left(\eta_{\mathcal{M}_{\text{off}}}(\mathcal{D}_{\text{off}})-\eta_{\mathcal{M}_{\text{on}}}(\pi)\right)$
$=-\frac{1}{1-\gamma}\,\mathbb{E}_{(s,a)\sim\rho_{\mathcal{M}_{\text{off}}}^{\mathcal{D}_{\text{off}}}}\left[\mathbb{E}_{s^{\prime}_{\text{off}}\sim p_{\mathcal{M}_{\text{off}}},\,a^{\prime}\sim\pi_{\mathcal{D}_{\text{off}}}}\left[Q^{\pi}_{\mathcal{M}_{\text{on}}}(s^{\prime}_{\text{off}},a^{\prime})\right]-\mathbb{E}_{s^{\prime}_{\text{on}}\sim p_{\mathcal{M}_{\text{on}}},\,a^{\prime}\sim\pi}\left[Q^{\pi}_{\mathcal{M}_{\text{on}}}(s^{\prime}_{\text{on}},a^{\prime})\right]\right]$
$=-\frac{1}{1-\gamma}\,\mathbb{E}_{(s,a)\sim\rho_{\mathcal{M}_{\text{off}}}^{\mathcal{D}_{\text{off}}}}\left[\right$.
$\quad\underbrace{\left(\mathbb{E}_{s^{\prime}_{\text{off}}\sim p_{\mathcal{M}_{\text{off}}},\,a^{\prime}\sim\pi_{\mathcal{D}_{\text{off}}}}\left[Q^{\pi}_{\mathcal{M}_{\text{on}}}(s^{\prime}_{\text{off}},a^{\prime})\right]-\mathbb{E}_{s^{\prime}_{\text{off}}\sim p_{\mathcal{M}_{\text{off}}},\,a^{\prime}\sim\pi}\left[Q^{\pi}_{\mathcal{M}_{\text{on}}}(s^{\prime}_{\text{off}},a^{\prime})\right]\right)}_{\text{(c)}}$
@@
$\mathbb{E}_{s^{\prime}_{\text{off}}\sim p_{\mathcal{M}_{\text{off}}},\,a^{\prime}\sim\pi_{\mathcal{D}_{\text{off}}}}\left[Q^{\pi}_{\mathcal{M}_{\text{on}}}(s^{\prime}_{\text{off}},a^{\prime})\right]-\mathbb{E}_{s^{\prime}_{\text{off}}\sim p_{\mathcal{M}_{\text{off}}},\,a^{\prime}\sim\pi}\left[Q^{\pi}_{\mathcal{M}_{\text{on}}}(s^{\prime}_{\text{off}},a^{\prime})\right]$
$\leq\mathbb{E}_{s^{\prime}_{\text{off}}\sim p_{\mathcal{M}_{\text{off}}}}\left[\sum_{a^{\prime}\in\mathcal{A}}\left|\mathcal{D}_{\text{off}}(a^{\prime}\mid s^{\prime}_{\text{off}})-\pi(a^{\prime}\mid s^{\prime}_{\text{off}})\right|\cdot\left|Q^{\pi}_{\mathcal{M}_{\text{on}}}(s^{\prime}_{\text{off}},a^{\prime})\right|\right]$
@@
$\leq\frac{2r_{\max}}{1-\gamma}\,\mathbb{E}_{s^{\prime}_{\text{off}}\sim p_{\mathcal{M}_{\text{off}}}}\left[D_{\text{TV}}\left(\pi_{\mathcal{D}_{\text{off}}}(\cdot\mid s^{\prime}_{\text{off}})\,\|\,\pi(\cdot\mid s^{\prime}_{\text{off}})\right)\right],$
@@
where the last inequality comes from the fact that $|Q^{\pi}_{\mathcal{M}_{\text{on}}}(s^{\prime}_{\text{off}},a^{\prime})|\leq\frac{r_{\rm max}}{1-\gamma}$ and the definition of TV distance.
$(d)$ $=\mathbb{E}_{s^{\prime}\sim p_{\mathcal{M}_{\text{off}}},\,a^{\prime}\sim\pi}\left[Q^{\pi}_{\mathcal{M}_{\text{on}}}(s^{\prime},a^{\prime})\right]-\mathbb{E}_{s^{\prime}\sim p_{\mathcal{M}_{\text{on}}},\,a^{\prime}\sim\pi}\left[Q^{\pi}_{\mathcal{M}_{\text{on}}}(s^{\prime},a^{\prime})\right]$
$=\int_{\mathcal{S}}\left(p_{\mathcal{M}_{\text{off}}}(s^{\prime}\mid s,a)-p_{\mathcal{M}_{\text{on}}}(s^{\prime}\mid s,a)\right)\left(\sum_{a^{\prime}}\pi(a^{\prime}\mid s^{\prime})Q^{\pi}_{\mathcal{M}_{\text{on}}}(s^{\prime},a^{\prime})\right)ds^{\prime}$
$\leq\frac{r_{\max}}{1-\gamma}\int_{\mathcal{S}}\left|p_{\mathcal{M}_{\text{off}}}(s^{\prime}\mid s,a)-p_{\mathcal{M}_{\text{on}}}(s^{\prime}\mid s,a)\right|ds^{\prime}$
$=\frac{2r_{\max}}{1-\gamma}\left[D_{\text{TV}}\left(p_{\mathcal{M}_{\text{off}}}(\cdot\mid s,a)\,\|\,p_{\mathcal{M}_{\text{on}}}(\cdot\mid s,a)\right)\right],$
where the inequality comes from the fact that $Q^{\pi}_{\mathcal{M}_{\text{on}}}(s^{\prime},a^{\prime})$ weighted by probability is bounded by $\frac{r_{\rm max}}{1-\gamma}$, the upper bound for all Q-values.
Combine the bounds for (c) and (d), we obtain a bound for the (a) term:
@@
$\eta_{\mathcal{M}_{\text{on}}}(\pi)-\eta_{\mathcal{M}_{\text{off}}}(\pi_{\mathcal{D}_{\text{off}}})$ $\geq-\frac{2r_{\max}}{(1-\gamma)^{2}}\,\mathbb{E}_{(s,a)\sim\rho^{\pi_{\mathcal{D}_{\text{off}}}}_{\mathcal{M}_{\text{off}}},\,s^{\prime}\sim p_{\mathcal{M}_{\text{off}}}}\left[D_{\text{TV}}\left(\pi_{\mathcal{D}_{\text{off}}}(\cdot\mid s^{\prime})\,\|\,\pi(\cdot\mid s^{\prime})\right)\right]$
$\quad-\frac{2r_{\max}}{(1-\gamma)^{2}}\,\mathbb{E}_{(s,a)\sim\rho^{\pi_{\mathcal{D}_{\text{off}}}}_{\mathcal{M}_{\text{off}}}}\left[D_{\text{TV}}\left(p_{\mathcal{M}_{\text{off}}}(\cdot\mid s,a)\,\|\,p_{\mathcal{M}_{\text{on}}}(\cdot\mid s,a)\right)\right].$
Now we try to bound term (b).
$\eta_{\mathcal{M}_{\text{off}}}(\pi_{\mathcal{D}_{\text{off}}})-\eta_{\mathcal{M}_{\text{off}}}(\pi)$
$=\frac{1}{1-\gamma}\,\mathbb{E}_{(s,a)\sim\rho^{\pi_{\mathcal{D}_{\text{off}}}}_{\mathcal{M}_{\text{off}}},\,s^{\prime}\sim p_{\mathcal{M}_{\text{off}}}(\cdot\mid s,a)}\left[\mathbb{E}_{a^{\prime}\sim\pi_{\mathcal{D}_{\text{off}}}}\left[Q^{\pi}_{\mathcal{M}_{\text{off}}}(s^{\prime},a^{\prime})\right]-\mathbb{E}_{a^{\prime}\sim\pi}\left[Q^{\pi}_{\mathcal{M}_{\text{off}}}(s^{\prime},a^{\prime})\right]\right]$
$\geq-\frac{1}{1-\gamma}\,\mathbb{E}_{(s,a)\sim\rho^{\pi_{\mathcal{D}_{\text{off}}}}_{\mathcal{M}_{\text{off}}},\,s^{\prime}\sim p_{\mathcal{M}_{\text{off}}}(\cdot\mid s,a)}\left|\mathbb{E}_{a^{\prime}\sim\pi_{\mathcal{D}_{\text{off}}}}\left[Q^{\pi}_{\mathcal{M}_{\text{off}}}(s^{\prime},a^{\prime})\right]-\mathbb{E}_{a^{\prime}\sim\pi}\left[Q^{\pi}_{\mathcal{M}_{\text{off}}}(s^{\prime},a^{\prime})\right]\right|$
$\geq-\frac{1}{1-\gamma}\,\mathbb{E}_{(s,a)\sim\rho^{\pi_{\mathcal{D}_{\text{off}}}}_{\mathcal{M}_{\text{off}}},\,s^{\prime}\sim p_{\mathcal{M}_{\text{off}}}(\cdot\mid s,a)}\left|\sum_{a^{\prime}\in\mathcal{A}}\left(\pi_{\mathcal{D}_{\text{off}}}(a^{\prime}\mid s^{\prime})-\pi(a^{\prime}\mid s^{\prime})\right)Q^{\pi}_{\mathcal{M}_{\text{off}}}(s^{\prime},a^{\prime})\right|$
$\geq-\frac{r_{\max}}{(1-\gamma)^{2}}\,\mathbb{E}_{(s,a)\sim\rho^{\pi_{\mathcal{D}_{\text{off}}}}_{\mathcal{M}_{\text{off}}},\,s^{\prime}\sim p_{\mathcal{M}_{\text{off}}}(\cdot\mid s,a)}\left|\sum_{a^{\prime}\in\mathcal{A}}\left(\pi_{\mathcal{D}_{\text{off}}}(a^{\prime}\mid s^{\prime})-\pi(a^{\prime}\mid s^{\prime})\right)\right|$
$=-\frac{2r_{\max}}{(1-\gamma)^{2}}\,\mathbb{E}_{(s,a)\sim\rho^{\pi_{\mathcal{D}_{\text{off}}}}_{\mathcal{M}_{\text{off}}},\,s^{\prime}\sim p_{\mathcal{M}_{\text{off}}}(\cdot\mid s,a)}\left[D_{\text{TV}}\left(\pi_{\mathcal{D}_{\text{off}}}(\cdot\mid s^{\prime})\,\|\,\pi(\cdot\mid s^{\prime})\right)\right],$
where the first equality is a direct application of Lemma  F.2. Combine the bound for term (a) and (b), we conclude that
$\eta_{\mathcal{M}_{\text{on}}}(\pi)-\eta_{\mathcal{M}_{\text{off}}}(\pi)\geq\;$ $-2C_{1}\,\mathbb{E}_{(s,a)\sim\rho_{\mathcal{M}}^{\pi_{\mathcal{D}_{\text{off}}}},s^{\prime}\sim p_{\text{off}}}\left[D_{\mathrm{TV}}(\pi(\cdot|s^{\prime})\parallel\pi_{\mathcal{D}_{\text{off}}}(\cdot|s^{\prime}))\right]$
$-C_{1}\,\mathbb{E}_{(s,a)\sim\rho_{\mathcal{M}}^{\pi_{\mathcal{D}_{\text{off}}}}}\left[D_{\mathrm{TV}}(p_{\text{on}}(\cdot|s,a)\parallel p_{\text{off}}(\cdot|s,a))\right],$
where $C_{1}=\frac{2r_{\max}}{(1-\gamma)^{2}}$. ∎
Appendix G Proof of Theorem  3.1
Definition G.1.